Warwicks Co, a womens clothing store, purchased $22,000 of merchandise from a supplier on account, terms FOB destination, 1/10, n/30. Warwicks returned $3,100 of the merchandise, recelving a credit memo. a. Journalize Warwicks Co.s entry to record the purchase. Inventory 21,780 Accounts Payable 21,780 Check My Work a. A purchase increases Merchandise Inventory and Accounts Payable Learning Objective 2 b. Journalize Warwicks Co.s entry to record the merchandise return Accounts Payable 3,069 Inventory 3,069 1 Check My Work b. A return reduces Merchandise Inventory and Accounts Payable in the purchasers record Leaming Objective 2
media%2F0eb%2F0eb422f6-1025-440c-83d7-1a
Answer #1

Date

Particulars

Debit

Credit

Accounts Payable

18900

          Cash

18900

          Merchandise Inventory

blank

blank

If the payment is made within ten days, discount is allowed. The discount amount is calculated as

Total inventory amount = 22000 – 3100 = 18900

Discount at 1/10 n/30 = 189

If the payment is made within ten days, the accounts payable is 18900-189 = 18711

If the payment is made beyond the ten days, the total amount is needed to be paid without discount reduction.